# Risk-Aware Adaptive Control Barrier Functions for Safe Control of Nonlinear Systems under Stochastic Uncertainty

### v3

@@ I Introduction
Constrained optimal control problems with safety requirements are fundamental in emerging safety-critical autonomous and cyber-physical systems, which are often modeled using stochastic systems. These systems must operate safely in unpredictable environments while interacting with humans, whose behavior is difficult to predict. To address risk, systems science and control engineering offer a rich history of methods for handling uncertainty, broadly categorized into worst-case, risk-neutral, and risk-aware approaches [1]. The risk-aware paradigm offers a balanced approach between worst-case and risk-neutral methods. While the worst-case paradigm ensures safety by considering the most severe violations [2], it often leads to overly conservative and impractical solutions. On the other hand, the risk-neutral paradigm [3] optimizes average performance but fails to account for rare but critical failures. Risk-aware methods address these limitations by considering outcomes beyond the average while avoiding excessive conservatism [4], leading to more practical and adaptive decision-making in safety-critical systems.
@@
Recent studies have demonstrated that stabilizing an affine control system while ensuring safety constraints and control limitations can be achieved by unifying Control Barrier Functions (CBFs) with Control Lyapunov Functions (CLFs), forming a sequence of single-step optimization programs [5, 6]. When using a quadratic cost, these optimizations become quadratic programs (QP), enabling real-time implementation [5]. Extensions of safety-critical control with CBFs have introduced formulations for high relative degree constraints and adaptive control [7, 8, 9, 10, 11]. To address safety in discrete-time systems, discrete-time CBFs (DCBFs) were introduced in [12], while discrete-time high-order CBFs (DHOCBFs) were later proposed in [13] to handle high relative degree constraints. All the CBFs mentioned above are for deterministic systems.
@@
Various approaches have been proposed in the literature to enforce safety in stochastic systems using CBFs. One approach to incorporating robust control in CBFs is to enforce safety by considering worst-case bounds on random variables and uncertainties [14, 15]. However, as previously mentioned, this often results in overly conservative behavior, keeping the system state far from the safe set boundary. A second approach ensures probabilistic satisfaction of safety constraints, meaning the safety condition holds with a probability exceeding a given threshold. This has been incorporated into the CBF framework using stochastic (Itô) differential equations [16, 17] or chance constraints [18]. Another approach uses extended class $\kappa$ functions instead of class $\kappa$ functions when defining a CBF [19, 20], ensuring that the system state is always stabilized to the safe set whenever the safety constraint is violated. However, these methods are unsuitable for scenarios where a single failure could result in catastrophic consequences. The recently proposed risk-aware DCBF approach [21, 22] provides a middle ground between worst-case robustness and probabilistic safety constraint satisfaction by explicitly considering both the likelihood and severity of undesirable outcomes. However, this approach is applicable exclusively to safety constraints that possess a relative degree of one. This means that if a safety constraint has a high relative degree, the uncertainties will propagate dynamically over multiple steps corresponding to the relative degree, making them difficult to compensate for with low-order constraints and potentially compromising safety. Moreover, the optimization problems in the aforementioned studies often become infeasible due to conflicts between stochastic safety constraints and control bounds. Additionally, these works do not address improving feasibility under such constraints.
@@ II-A Discrete-Time High-Order CBFs
Definition 1 (Relative degree [23]).
@@
Definition 2 (DHOCBF [13]).
@@
Theorem 1 (Safety Guarantee [13]).
@@ II-B Coherent Risk Measure
A fundamental property of dynamic risk measures is their consistency over time [24]. This means that if a risk evaluation at some future time $t_{\tau}$ considers $h$ to be as good as $h^{\prime}$, and if $h$ and $h^{\prime}$ are identical between times $t_{k}$ and $t_{\tau}$ ($k<\tau$), then from the perspective of time $t_{k}$, $h$ should not be considered worse than $h^{\prime}$. For a time-consistent risk measure, the one-step conditional risk measure $\rho_{t}$ is defined as a mapping:
@@
However, VaR is numerically unstable for non-normal stochastic variables, and optimization problems involving VaR become intractable in high-dimensional settings. Moreover, VaR ignores the magnitude of values below the $\beta$-quantile [25]. In contrast, CVaR addresses these issues by considering the expected loss in the $\beta$-tail. The standard CVaR with confidence level $\beta$ is given by:
@@
An alternative formulation introduced in [25] is:
@@
This formulation is used in this paper instead of (11) because it is more tractable for optimization and ensures time consistency in sequential decision-making. The standard CVaR formula (11) is non-convex and difficult to optimize, whereas Eq. (12) reformulates CVaR in a convex and computationally efficient way. Additionally, the standard CVaR definition does not extend naturally to multi-stage problems, making it unsuitable for risk-aware control applications. Equation (12) provides a flexible framework that integrates well with dynamic models, making it more practical for real-world safety-critical systems. A confidence level $\beta\to 1$ corresponds to a risk-neutral case, where $\text{CVaR}_{1}(h)=\mathbb{E}[h]$. Conversely, $\beta\to 0$ results in a risk-averse measure, where $\text{CVaR}_{0}(h)=\text{VaR}_{0}(h)=\operatorname{ess}\inf(h)$ [26].
@@ III Problem Formulation and Approach
$\begin{split}J(\mathbf{u}_{t},\mathbf{x}_{t+1})=\mathbf{u}_{t}^{T}Q\mathbf{u}_{t}+(\mathbf{x}_{t+1}-\mathbf{x}_{e})^{T}R(\mathbf{x}_{t+1}-\mathbf{x}_{e}),\end{split}$ (14)
@@ IV-A Enhancing Optimization Feasibility with DAVCBFs
Motivated by the auxiliary-variable adaptive CBFs in [9], given a function $h:\mathbb{R}^{n}\to\mathbb{R}$ with relative degree $m$ for system (1), and a time-varying vector of auxiliary variables $\bm{a}_{t}\coloneqq[a_{1,t},\dots,a_{m,t}]^{T}$ with positive components, we place one auxiliary variable in front of each function $\psi_{i}(\mathbf{x})$ in (3). If each auxiliary variable is ensured to always be positive as $a_{i,t}>0,\quad i\in\{1,\dots,m\}$, then multiplying the auxiliary variable $a_{i,t}$ in front of $\psi_{i-1}(\mathbf{x})$ is equivalent to enforcing $\psi_{i-1}(\mathbf{x})\geq 0$. Therefore, the forward invariance of the intersection of a sequence of sets $\mathcal{C}_{i}$ defined by (4) can still be rendered with auxiliary variables. To ensure that $a_{i,t}$ remains positive, we define a DHOCBF for each $a_{i,t}$ and introduce auxiliary systems containing auxiliary states $\bm{\pi}_{i,t}$ and inputs $\nu_{i,t}$. Through these systems, we extend each DHOCBF to the desired relative degree, similar to how $h(\mathbf{x})$ has relative degree $m$ with respect to the dynamics (1). Consider $m$ auxiliary systems in the form:
@@ IV-D Optimal Control with RACBFs
$\begin{split}J(\mathbf{u}_{t},\mathbf{x}_{t+1},\bm{\nu}_{t},\sigma_{i,t},\zeta_{t})=\mathbf{u}_{t}^{T}Q\mathbf{u}_{t}+(\mathbf{x}_{t+1}-\mathbf{x}_{e})^{T}R(\mathbf{x}_{t+1}-\mathbf{x}_{e})\\ +(\bm{\nu}_{t}-\mathbf{a}_{w})^{T}W(\bm{\nu}_{t}-\mathbf{a}_{w})+P(\zeta_{t}+\frac{1}{\beta|\mathcal{W}|}\sum_{i=1}^{|\mathcal{W}|}\sigma_{i,t}).\end{split}$ (30)
@@
$\begin{split}-(\zeta_{t}+\frac{1}{\beta|\mathcal{W}|}\sum_{i=1}^{|\mathcal{W}|}\sigma_{i,t})\geq 0,\sigma_{i,t}\geq 0,\\ \sigma_{i,t}\geq-\rho^{i}(\psi_{m}(\mathbf{x}_{t+1},\bm{\Pi}_{t+1}))-\zeta_{t},i\in\{1,...,|\mathcal{W}|\}.\end{split}$ (31)
@@ V CASE STUDY AND SIMULATIONS
In this section, we present numerical results to validate our proposed approach using a unicycle robot modeled as a nonlinear discrete-time system. All computations and simulations were conducted in MATLAB, where the nonlinear optimization problem was solved using IPOPT [27] with the modeling language Yalmip [28]. Consider a stochastic discrete-time unicycle model in the form
@@
$\begin{split}\min_{\mathbf{u}_{t},\mathbf{x}_{t+1},\nu_{1,t},\sigma_{i,t},\zeta_{t}}[\mathbf{u}_{t}^{T}Q\mathbf{u}_{t}+(\mathbf{x}_{t+1}-\mathbf{x}_{e})^{T}R(\mathbf{x}_{t+1}-\mathbf{x}_{e})\\ +W(\nu_{1,t}-a_{1,w})^{2}+P(\zeta_{t}+\frac{1}{\beta|\mathcal{W}|}\sum_{i=1}^{|\mathcal{W}|}\sigma_{i,t})].\end{split}$ (38)

### v5

@@ I Introduction
Constrained optimal control problems with safety requirements are fundamental in emerging safety-critical autonomous and cyber-physical systems, which are often modeled using stochastic systems. These systems must operate safely in unpredictable environments while interacting with humans, whose behavior is difficult to predict. To address risk, systems science and control engineering offer a rich history of methods for handling uncertainty, broadly categorized into worst-case, risk-neutral, and risk-aware approaches [wang2022risk]. The risk-aware paradigm offers a balanced approach between worst-case and risk-neutral methods. While the worst-case paradigm ensures safety by considering the most severe violations [zhou1998essentials], it often leads to overly conservative and impractical solutions. On the other hand, the risk-neutral paradigm [bertsekas1996stochastic] optimizes average performance but fails to account for rare but critical failures. Risk-aware methods address these limitations by considering outcomes beyond the average while avoiding excessive conservatism [howard1972risk], leading to more practical and adaptive decision-making in safety-critical systems.
@@
Recent studies have demonstrated that stabilizing an affine control system while ensuring safety constraints and control limitations can be achieved by unifying Control Barrier Functions (CBFs) with Control Lyapunov Functions (CLFs), forming a sequence of single-step optimization programs [ames2014control, ames2016control]. When using a quadratic cost, these optimizations become quadratic programs (QP), enabling real-time implementation [ames2014control]. Extensions of safety-critical control with CBFs have introduced formulations for high relative degree constraints and adaptive control [nguyen2016exponential, xiao2021high, liu2023auxiliary, liu2024auxiliary, liu2025auxiliary]. To address safety in discrete-time systems, discrete-time CBFs (DCBFs) were introduced in [agrawal2017discrete], while discrete-time high-order CBFs (DHOCBFs) were later proposed in [xiong2022discrete] to handle high relative degree constraints. All the CBFs mentioned above are for deterministic systems.
@@
Various approaches have been proposed in the literature to enforce safety in stochastic systems using CBFs. One approach to incorporating robust control in CBFs is to enforce safety by considering worst-case bounds on random variables and uncertainties [lindemann2019control, cohen2022high]. However, as previously mentioned, this often results in overly conservative behavior, keeping the system state far from the safe set boundary. A second approach ensures probabilistic satisfaction of safety constraints, meaning the safety condition holds with a probability exceeding a given threshold. This has been incorporated into the CBF framework using stochastic (Itô) differential equations [clark2021control, santoyo2021barrier] or chance constraints [wang2021chance]. Another approach uses extended class $\kappa$ functions instead of class $\kappa$ functions when defining a CBF [jankovic2018robust, tan2021high], ensuring that the system state is always stabilized to the safe set whenever the safety constraint is violated. However, these methods are unsuitable for scenarios where a single failure could result in catastrophic consequences. The recently proposed risk-aware DCBF approach [singletary2022safe, akella2024risk] provides a middle ground between worst-case robustness and probabilistic safety constraint satisfaction by explicitly considering both the likelihood and severity of undesirable outcomes. However, this approach is applicable exclusively to safety constraints that possess a relative degree of one. This means that if a safety constraint has a high relative degree, the uncertainties will propagate dynamically over multiple steps corresponding to the relative degree, making them difficult to compensate for with low-order constraints and potentially compromising safety. Moreover, the optimization problems in the aforementioned studies often become infeasible due to conflicts between stochastic safety constraints and control bounds. Additionally, these works do not address improving feasibility under such constraints.
@@ II-A Discrete-Time High-Order CBFs
Definition 1 (Relative degree [sun2003initial]).
@@
Definition 2 (DHOCBF [xiong2022discrete]).
@@
Theorem 1 (Safety Guarantee [xiong2022discrete]).
@@ II-B Coherent Risk Measure
A fundamental property of dynamic risk measures is their consistency over time [ruszczynski2010risk]. This means that if a risk evaluation at some future time $t_{\tau}$ considers $h$ to be as good as $h^{\prime}$, and if $h$ and $h^{\prime}$ are identical between times $t_{k}$ and $t_{\tau}$ ($k<\tau$), then from the perspective of time $t_{k}$, $h$ should not be considered worse than $h^{\prime}$. For a time-consistent risk measure, the one-step conditional risk measure $\rho_{t}$ is defined as a mapping:
@@
However, VaR is numerically unstable for non-normal stochastic variables, and optimization problems involving VaR become intractable in high-dimensional settings. Moreover, VaR ignores the magnitude of values below the $\beta$-quantile [rockafellar2000optimization]. In contrast, CVaR addresses these issues by considering the expected loss in the $\beta$-tail. The standard CVaR with confidence level $\beta$ is given by:
@@
An alternative formulation introduced in [rockafellar2000optimization] is:
@@
This formulation is used in this paper instead of (11) because it is more tractable for optimization and ensures time consistency in sequential decision-making. The standard CVaR formula (11) is non-convex and difficult to optimize, whereas Eq. (12) reformulates CVaR in a convex and computationally efficient way. Additionally, the standard CVaR definition does not extend naturally to multi-stage problems, making it unsuitable for risk-aware control applications. Equation (12) provides a flexible framework that integrates well with dynamic models, making it more practical for real-world safety-critical systems. A confidence level $\beta\to 1$ corresponds to a risk-neutral case, where $\text{CVaR}_{1}(h)=\mathbb{E}[h]$. Conversely, $\beta\to 0$ results in a risk-averse measure, where $\text{CVaR}_{0}(h)=\text{VaR}_{0}(h)=\operatorname{ess}\inf(h)$ [rockafellar2002conditional].
@@ III Problem Formulation and Approach
$\begin{split}J(\mathbf{u}_{t},\mathbf{x}(\mathbf{u}_{t}))=\mathbf{u}_{t}^{T}Q\mathbf{u}_{t}+(\mathbf{x}(\mathbf{u}_{t})-\mathbf{x}_{e})^{T}R(\mathbf{x}(\mathbf{u}_{t})-\mathbf{x}_{e}),\end{split}$ (14)
@@ IV-A Enhancing Optimization Feasibility with DAVCBFs
Motivated by the auxiliary-variable adaptive CBFs in [liu2023auxiliary], given a function $h:\mathbb{R}^{n}\to\mathbb{R}$ with relative degree $m$ for system (1), and a time-varying vector of auxiliary variables $\bm{a}_{t}\coloneqq[a_{1,t},\dots,a_{m,t}]^{T}$ with positive components, we place one auxiliary variable in front of each function $\psi_{i}(\mathbf{x})$ in (3). If each auxiliary variable is ensured to always be positive as $a_{i,t}>0,\quad i\in\{1,\dots,m\}$, then multiplying the auxiliary variable $a_{i,t}$ in front of $\psi_{i-1}(\mathbf{x})$ is equivalent to enforcing $\psi_{i-1}(\mathbf{x})\geq 0$. Therefore, the forward invariance of the intersection of a sequence of sets $\mathcal{C}_{i}$ defined by (4) can still be rendered with auxiliary variables. To ensure that $a_{i,t}$ remains positive, we define a DHOCBF for each $a_{i,t}$ and introduce auxiliary systems containing auxiliary states $\bm{\pi}_{i,t}$ and inputs $\nu_{i,t}$. Through these systems, we extend each DHOCBF to the desired relative degree, similar to how $h(\mathbf{x})$ has relative degree $m$ with respect to the dynamics (1). Consider $m$ auxiliary systems in the form:
@@ IV-D Optimal Control with RACBFs
$\begin{split}J(\mathbf{u}_{t},\mathbf{x}(\mathbf{u}_{t}),\bm{\nu}_{t},\sigma_{i,t},\zeta_{t})=\mathbf{u}_{t}^{T}Q\mathbf{u}_{t}+(\mathbf{x}(\mathbf{u}_{t})-\mathbf{x}_{e})^{T}R(\mathbf{x}(\mathbf{u}_{t})-\mathbf{x}_{e})\\ +(\bm{\nu}_{t}-\mathbf{a}_{w})^{T}W(\bm{\nu}_{t}-\mathbf{a}_{w})+P(\zeta_{t}+\frac{1}{\beta|\mathcal{W}|}\sum_{i=1}^{|\mathcal{W}|}\sigma_{i,t}).\end{split}$ (30)
@@
$\begin{split}-(\zeta_{t}+\frac{1}{\beta|\mathcal{W}|}\sum_{i=1}^{|\mathcal{W}|}\sigma_{i,t})\geq 0,\sigma_{i,t}\geq 0,\\ \sigma_{i,t}\geq-\rho^{i}(\psi_{m}(\mathbf{x}_{t},\bm{\Pi}_{t}))-\zeta_{t},i\in\{1,...,|\mathcal{W}|\}.\end{split}$ (31)
@@ V CASE STUDY AND SIMULATIONS
In this section, we present numerical results to validate our proposed approach using a unicycle robot modeled as a nonlinear discrete-time system. All computations and simulations were conducted in MATLAB, where the nonlinear optimization problem was solved using IPOPT [biegler2009large] with the modeling language Yalmip [lofberg2004yalmip]. Consider a stochastic discrete-time unicycle model in the form
@@
$\begin{split}\min_{\mathbf{u}_{t},\mathbf{x}(\mathbf{u}_{t}),\nu_{1,t},\sigma_{i,t},\zeta_{t}}[\mathbf{u}_{t}^{T}Q\mathbf{u}_{t}+(\mathbf{x}(\mathbf{u}_{t})-\mathbf{x}_{e})^{T}R(\mathbf{x}(\mathbf{u}_{t})-\mathbf{x}_{e})\\ +W(\nu_{1,t}-a_{1,w})^{2}+P(\zeta_{t}+\frac{1}{\beta|\mathcal{W}|}\sum_{i=1}^{|\mathcal{W}|}\sigma_{i,t})].\end{split}$ (38)